OpenStudy (anonymous):

The table describe an inverse variation relationship between x & y. Write the equation. x: 3, 5, 9, 10, 15, 25, 45, 90 y: 90, 45, 25, 22.5, 15, 9, 5, 3 Produce Form: Y=Form:

OpenStudy (anonymous):

HELP!

OpenStudy (anonymous):

@lalaly - Please help me with this question!

OpenStudy (lalaly):

are u sure thats a 3? above the 90?

OpenStudy (anonymous):

Yes! Im really confusing

OpenStudy (lalaly):

i dont know about the first one i think its supposed to be 2.5 anyways the equation is y=k/x and xy=k for all the values except 90 and 3... k= 225 so the equation is y=225/x

OpenStudy (anonymous):

okay Thanks :)